# Nearsightedness in Materials with Indirect Band Gap

### v1

@@ 6. Proofs of the Main Results
$\leq C\Big{(}\sum_{\begin{subarray}{c}m,n\in B_{R}\end{subarray}}F(\ell,m,n,k)+\sum_{{m\notin B_{R}~{}\text{or}~{}n\notin B_{R}}}F(\ell,m,n,k)\Big{)}$
$=:C\big{(}T_{1}+T_{2}\big{)},$ (6.26)
where we define
$F(\ell,m,n,k):=\sum_{1\leq b_{1},c_{1}\leq N_{\mathrm{b}}}\Big{|}\big{(}z-{\mathcal{H}}(u)\big{)}^{-1}_{\ell m,ab_{1}}\Big{|}\cdot\Big{|}\big{[}{\mathcal{H}}(u)-{\mathcal{H}}^{\mathrm{ref}}\big{]}_{mn,b_{1}c_{1}}\Big{|}\cdot\Big{|}(z-{\mathcal{H}}^{\mathrm{ref}})^{-1}_{nk,c_{1}b}\Big{|}.$
Here, we choose a sufficiently large $R$ in (6.2) so that (6.15) still holds. Combining the regularity of Hamiltonians with Combes–Thomas estimate (3.4) and its improved version (6.24), we have
$T_{1}$ $=\sum_{m\in B_{R}~{}\text{or}~{}n\in B_{R}}F(\ell,m,n,k)$
$\leq\tilde{C}(u)\sum_{m\in B_{R}~{}\text{or}~{}n\in B_{R}}e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell m}}e^{-\gamma_{0}\mathfrak{m}r_{mn}}e^{-\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref})r_{mk}}$
$\leq\tilde{C}(u)\sum_{m\in B_{R}}e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell m}}e^{-\frac{1}{2}\min\{\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref}),\gamma_{0}\mathfrak{m}\}r_{nk}}\leq\tilde{C}(u)\sum_{m\in B_{R}}e^{-4\eta_{-}(r_{\ell m}+r_{mk})}$
$\leq\tilde{C}(u)\Big{(}e^{-2\eta_{-}\min_{m\in B_{R}}(r_{\ell m}+r_{mk})}\sum_{m\in\Lambda}e^{-2\eta_{-}(r_{\ell m}+r_{mk})}\Big{)}$
$\leq\tilde{C}(u)e^{-2\eta_{-}\min_{m\in B_{R}}(r_{\ell m}+r_{mk})}e^{-\eta_{-}r_{\ell k}}$ (6.27)
where we denote
$\eta_{-}:=\frac{1}{8}\min\{\gamma_{\rm{CT}}(\tfrac{1}{2}\mathfrak{d}^{\mathrm{ref}}),\gamma_{0}\mathfrak{m}\}=c_{0}\min\{h_{0},\gamma_{0}^{d}\mathfrak{d}^{\mathrm{ref}}\},$ (6.28)
with some $c_{0}>0$ depends only on $\mathfrak{m},d,h_{0},\gamma_{0}$. Then, using (6.15) and Hölder’s inequality yields
$T_{2}=\sum_{m,n\in\Lambda\setminus B_{R}}F(\ell,m,n,k)$
$\leq\tilde{C}(u)\sum_{m,n\in\Lambda\setminus B_{R}}e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell m}}\Big{(}e^{-\frac{\sqrt{3}}{2}\gamma_{0}\mathfrak{m}r_{mn}}|u(m)-u(n)|\Big{)}e^{-\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref})r_{nk}}$
$\leq\tilde{C}(u)\Big{(}\sum_{m,n\in\Lambda\setminus B_{R}}e^{-\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell m}}e^{-(\sqrt{3}\gamma_{0}\mathfrak{m}-2\Upsilon)r_{mn}}e^{-2\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref})r_{nk}}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{m,n\in\Lambda\setminus B_{R}}e^{-2\Upsilon r_{mn}}|u(m)-u(n)|^{2}\Big{)}^{2}$
@@
$\leq\tilde{C}(u)\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R})}\Big{(}\sum_{m,n\in\Lambda\setminus B_{R}}e^{-\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell m}}e^{-\frac{\sqrt{3}}{2}\gamma_{0}\mathfrak{m}r_{mn}}e^{-2\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref})r_{nk}}\Big{)}^{\frac{1}{2}}$
$\leq\tilde{C}(u)\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R})}\sum_{m,n\in\Lambda\setminus B_{R}}e^{-4\eta_{-}(r_{\ell m}+r_{mn}+r_{nk})}\leq\tilde{C}(u)\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R})}\sum_{m\in\Lambda}e^{-2\eta_{-}(r_{\ell m}+r_{mk})}$
$\leq\tilde{C}(u)\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R})}e^{-\eta_{-}r_{\ell k}}.$ (6.29)
From Lemma C, we note that the prefactors in (6.2) and (6.2) are both independent of $\eta_{-}$ as $\mathsf{gap}_{-}$ is small. Moreover, by choosing $\mathscr{C}$ satisfying $\mathrm{dist}(\mathscr{C},\sigma({\mathcal{H}}^{\mathrm{ref}}))\geq\frac{1}{2}\mathsf{gap}_{-}$, we can take $\mathsf{gap}_{-}$ in $\eta_{-}$. Finally, substituting (6.2) and (6.2) into (6.2) draws the conclusion that
$\big{|}\rho(u)_{\ell k,ab}-\rho^{\mathrm{ref}}_{\ell k,ab}\big{|}\leq C(T_{1}+T_{2})$
@@
$\hskip 28.45274pt\leq\frac{C\|Du\|_{\ell^{2}_{\Upsilon}}}{\mathfrak{d}^{\mathrm{ref}}}\bigg{(}1+\frac{R_{1}^{d}e^{\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})R_{1}}}{(\mathfrak{d}^{\mathrm{ref}})^{d+3}\big{(}\mathfrak{d}(u)\big{)}^{d+1}}\bigg{)}\Big{(}\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R})}+e^{-2\eta_{-}\min\limits_{m\in B_{R}}[r_{\ell m}+r_{mk}]}\Big{)}e^{-\eta_{-}r_{\ell k}}$
$\hskip 28.45274pt\leq\frac{C\|Du\|_{\ell^{2}_{\Upsilon}}}{\mathsf{gap}_{-}}\bigg{(}1+\frac{R_{1}^{d}e^{4\eta_{-}R_{1}}}{\mathsf{gap}_{-}^{2d+4}}\bigg{)}\Big{(}\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R})}+e^{-2\eta_{-}\min\limits_{m\in B_{R}}[r_{\ell m}+r_{mk}]}\Big{)}e^{-\eta_{-}r_{\ell k}}$
$\hskip 28.45274pt:=C_{\ell k}(u)e^{-\eta_{-}r_{\ell k}}.$ (6.30)
Taking this estimate into (6.25) completes the proof. ∎
6.3. Proof of the Proposition 3.2
We first provide an estimate for resolvents in homogeneous system exhibiting super-algebraic decay, which helps clarify the dependence of the strong locality on decay rate.
Lemma 6.4. Suppose ${\mathcal{H}}^{\mathrm{ref}}$ satisfies (TB), and $z\in\mathscr{C}$ such that $\mathrm{dist}\left(z,\sigma\big{(}{\mathcal{H}}^{\mathrm{ref}}\big{)}\right)\geq\mathfrak{d}^{\mathrm{ref}}$ for some $\mathfrak{d}^{\mathrm{ref}}>0$. Then, for each $q\in\mathbb{Z}^{+}$, there exists a constant $C_{q}>0$ such that
$\left|\big{(}z-{\mathcal{H}}^{\mathrm{ref}}\big{)}^{-1}_{\ell k,ab}\right|$ $\leq\frac{C_{q}(\mathfrak{d}^{\mathrm{ref}})}{r_{\ell k}^{q}}e^{-\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref})r_{\ell k}},$ (6.31)
for all $\ell\neq k,\ell,k\in\Lambda$, and $1\leq a,b\leq N_{\mathrm{b}}$, where the constant is defined by
$C_{q}(\mathfrak{d}^{\mathrm{ref}}):=\begin{cases}cq!{\big{(}(\mathfrak{d}^{\mathrm{ref}})^{q}-1\big{)}}\big{/}{\big{(}(\mathfrak{d}^{\mathrm{ref}})^{q+1}(\mathfrak{d}^{\mathrm{ref}}-1)\big{)}},~{}~{}\mathfrak{d}^{\mathrm{ref}}\neq 1,\\ cq!(q/\mathfrak{d}^{\mathrm{ref}}),~{}~{}\mathfrak{d}^{\mathrm{ref}}=1,\end{cases}$ (6.32)
for some $c>0$ depending only on $h_{0},\gamma_{0},N_{\mathrm{b}},M$.
Proof.
Recall that in Appendix A, we proved there exists a decomposition for ${\mathcal{H}}^{\mathrm{ref}}$, denoted by ${\mathcal{H}}_{\boldsymbol{\xi}}\in\mathbb{R}^{MN_{\mathrm{b}}\times MN_{\mathrm{b}}}$, $\bm{\xi}\in\Gamma^{*}$. So the resolvent can be written as
[…]
By partial integration, we have
$(z-{\mathcal{H}}^{\mathrm{ref}})^{-1}_{\ell k,ab}$ $=\frac{i^{|\bm{\alpha}|}}{[\bm{r}_{\ell k}]^{\bm{\alpha}}}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.70335pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.59167pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.10834pt}}\!\int_{\Gamma^{\star}}\frac{\partial^{|\bm{\alpha}|}}{\partial\bm{\xi}^{\bm{\alpha}}}\left((z-{\mathcal{H}}_{\boldsymbol{\xi}})^{-1}_{\ell k,ab}\right)e^{i\bm{\xi}\cdot\bm{r}_{\ell k}}\mathrm{d}\bm{\xi},$ (6.34)
where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{d})\in\mathbb{N}^{d}$ is the multi-index with $|\bm{\alpha}|:=\sum_{i}\alpha_{i}$, and $[\bm{r}_{\ell k}]^{\bm{\alpha}}:=\prod_{i}[\bm{r}_{\ell k}]_{i}^{\alpha_{i}}$. The first derivative of $(z-{\mathcal{H}}_{\boldsymbol{\xi}})^{-1}$ is
$\frac{\partial}{\partial\bm{\xi}_{l}}\left((z-{\mathcal{H}}_{\boldsymbol{\xi}})^{-1}\right)$ $=(z-{\mathcal{H}}_{\boldsymbol{\xi}})^{-1}\frac{\partial{\mathcal{H}}_{\boldsymbol{\xi}}}{\partial\bm{\xi}_{l}}(z-{\mathcal{H}}_{\boldsymbol{\xi}})^{-1},\quad\forall 1\leq l\leq d.$ (6.35)
For $|\bm{\alpha}|=q\in\mathbb{N},q>1$, the higher order derivatives can be given by
$\frac{\partial^{|\bm{\alpha}|}}{\partial\bm{\xi}^{\bm{\alpha}}}\left((z-{\mathcal{H}}_{\boldsymbol{\xi}})^{-1}\right)$ $=\sum_{n=1}^{q}\sum_{\begin{subarray}{c}\bm{\beta}_{1},\dots,\bm{\beta}_{n}\in\mathbb{N}^{d}\setminus\{\bm{0}\}\colon\\ \bm{\alpha}=\bm{\beta}_{1}+\dots+\bm{\beta}_{n}\end{subarray}}\widehat{D}_{z,\bm{\xi}}^{(n)}\left(\frac{\partial^{|\bm{\beta}_{1}|}{\mathcal{H}}_{\boldsymbol{\xi}}}{\partial\bm{\xi}^{\bm{\beta}_{1}}},\dots,\frac{\partial^{|\bm{\beta}_{n}|}{\mathcal{H}}_{\boldsymbol{\xi}}}{\partial\bm{\xi}^{\bm{\beta}_{n}}}\right),$ (6.36)
where $\widehat{D}_{z,\bm{\xi}}^{(n)}:(\mathbb{R}^{MN_{\mathrm{b}}\times MN_{\mathrm{b}}})^{n}\to\mathbb{R}^{MN_{\mathrm{b}}\times MN_{\mathrm{b}}}$ is given by
$\widehat{D}_{z,\bm{\xi}}^{(n)}(X_{1},\dots,X_{n})=(z-{\mathcal{H}}_{\boldsymbol{\xi}})^{-1}\prod_{j=1}^{n}\left[X_{j}(z-{\mathcal{H}}_{\boldsymbol{\xi}})^{-1}\right].$ (6.37)
For all $\bm{\xi}\in\Gamma^{*}$, $\widehat{D}_{z,\bm{\xi}}^{(n)}$ is an analytic function as $\mathrm{dist}\left(z,\sigma\big{(}{\mathcal{H}}^{\mathrm{ref}}\big{)}\right)\geq\mathfrak{d}^{\mathrm{ref}}$. Thus, for all $|\bm{\zeta}|\leq\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref}),\bm{\zeta}\in\mathbb{R}^{d}$, we have
[…]
where we define $\big{[}{D}_{z}^{(n)}\big{]}_{\ell k,ab}:=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.70335pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.59167pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.10834pt}}\!\int_{\Gamma^{\star}}\big{[}\widehat{D}_{z,\bm{\xi}}^{(n)}\big{]}_{\ell_{0}k_{0},ab}e^{i\bm{\xi}\cdot\bm{r}_{\ell k}}\mathrm{d}\bm{\xi}.$ Moreover, by solving the eigenvalue problem (A.1), we have the eigenpairs $\big{(}\varepsilon_{n}(\bm{\xi}),u_{n}(\bm{\xi})\big{)}$ of ${\mathcal{H}}_{\boldsymbol{\xi}}$, where $\{u_{n}(\bm{\xi})\}_{n}$ are orthonormal and continues in $\bm{\xi}$, and then
$\big{|}(z-{\mathcal{H}}_{\boldsymbol{\xi}})^{-1}_{\ell_{0}k_{0},ab}\big{|}=\left|\sum_{n=1}^{MN_{\mathrm{b}}}\frac{[u_{n}(\bm{\xi})\otimes u_{n}(\bm{\xi})]_{\ell_{0}k_{0},ab}}{z-\varepsilon_{n}(\bm{\xi})}\right|\leq CMN_{b}(\mathfrak{d}^{\mathrm{ref}})^{-1},\quad\forall\bm{\xi}\in\Gamma^{*},$ (6.39)
for some $C>0$. From this and (6.37), we obtain
$\left|\widehat{D}_{z,\bm{\xi}+i\bm{\zeta}}^{(n)}(X_{1},\dots,X_{n})_{\ell_{0}k_{0},ab}\right|\leq C(\mathfrak{d}^{\mathrm{ref}})^{-(n+1)}\prod_{j=1}^{n}\|X_{j}\|_{\infty}$ (6.40)
with $X_{j}={\partial^{|\bm{\beta}_{j}|}{\mathcal{H}}_{\boldsymbol{\xi}}}/{\partial\bm{\xi}^{\bm{\beta}_{j}}},j=1,\dots,n,$ for some constant $C>0$ depending on $M,N_{b}$. Combining (6.38) with (6.40), in a similar manner as (6.10), we obtain
[…]
with a constant $C$ depending on $X_{j},j=1,\cdots,n$. Note that, for each $1\leq n\leq q$, there exits at most $d!n!\binom{n}{q}\lesssim q!$ terms of $\widehat{D}_{z,\bm{\xi}}^{(n)}$ for the second sum in (6.36). Using (6.34), (6.36) and (6.41), we arrive at
$\left|\big{(}z-{\mathcal{H}}^{\mathrm{ref}}\big{)}^{-1}_{\ell k,ab}\right|$ $\leq\frac{C}{r_{\ell k}^{q}}\sum_{n=1}^{q}\sum_{\begin{subarray}{c}\bm{\alpha}_{1},\dots,\bm{\alpha}_{n}\in\mathbb{N}^{d}\setminus\{\bm{0}\}\colon\\ |\bm{\alpha}_{1}|+\dots+|\bm{\alpha}_{n}|=q\end{subarray}}\left|\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.70335pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.59167pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.10834pt}}\!\int_{\Gamma^{\star}}\widehat{D}_{z,\bm{\xi}}^{(n)}\left(\frac{\partial^{\bm{\alpha}_{1}}{\mathcal{H}}_{\boldsymbol{\xi}}}{\partial\bm{\xi}_{\bm{\alpha}_{1}}},\cdots,\frac{\partial^{\bm{\alpha}_{n}}{\mathcal{H}}_{\boldsymbol{\xi}}}{\partial\bm{\xi}_{\bm{\alpha}_{n}}}\right)_{\ell_{0}k_{0},ab}e^{i\bm{\xi}\cdot\bm{r}_{\ell k}}\mathrm{d}\bm{\xi}\right|$
$\leq\frac{C}{r_{\ell k}^{q}}\sum_{n=1}^{q}\sum_{\begin{subarray}{c}\bm{\alpha}_{1},\dots,\bm{\alpha}_{n}\in\mathbb{N}^{d}\setminus\{\bm{0}\}\colon\\ |\bm{\alpha}_{1}|+\dots+|\bm{\alpha}_{n}|=q\end{subarray}}{(\mathfrak{d}^{\mathrm{ref}})^{-(n+1)}}e^{-\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref})r_{\ell k}}$
$\leq\frac{Cq!}{r_{\ell k}^{q}}\Big{(}{(\mathfrak{d}^{\mathrm{ref}})^{-2}}+{(\mathfrak{d}^{\mathrm{ref}})^{-3}}+\dots+{(\mathfrak{d}^{\mathrm{ref}})^{-(q+1)}}\Big{)}e^{-\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref})r_{\ell k}}$
$\leq\frac{C_{q}(\mathfrak{d}^{\mathrm{ref}})}{r_{\ell k}^{q}}e^{-\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref})r_{\ell k}},$ (6.42)
which completes the proof. ∎
Remark 6.5. The above result can be extended to ${\mathcal{H}}(u)$ using a similar argument as in Lemma 6.2 by replacing the estimate for $|(z-{\mathcal{H}}^{\mathrm{ref}})^{-1}_{\ell k,ab}|$ with (6.31), which yields
$\left|\big{(}z-{\mathcal{H}}(u)\big{)}^{-1}_{\ell k,ab}\right|\leq\frac{C_{q}(\mathfrak{d}^{\mathrm{ref}})}{r_{\ell k}^{q}}e^{-\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell k}}+C(u)e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})(|\ell|+|k|)},$ (6.43)
for all $\ell\neq k,\ell,k\in\Lambda$, and $1\leq a,b\leq N_{\mathrm{b}}$, where the constant $C(u)$ depends on $\mathfrak{d}(u)$ and $\|Du\|_{\ell^{2}_{\Upsilon}}$. ∎
We are now in a position to prove Proposition 3.2:
Proof of Proposition 3.2.
For the inhomogeneous case, by using (6.43), we have

### v2

@@ 6.2. Proof of Section 3
$\leq\sum_{mn}\oint_{\mathscr{C}}\Big|(z-{\mathcal{H}}(u))^{-1}_{\ell m}(P_{\delta}(u)+P_{\rm loc}(u))_{mn}(z-{\mathcal{H}}^{\rm ref})^{-1}_{nk}\Big|\frac{{\,{\rm d}}z}{2\pi i}$
$\leq C\sum_{mn}\left|(z-{\mathcal{H}}(u))^{-1}_{\ell m}\right|\left(\left|P_{\delta}(u)_{mn}\right|+\left|P_{\rm loc}(u))_{mn}\right|\right)\left|(z-{\mathcal{H}}^{\rm ref})^{-1}_{nk}\right|$
(6.27) $\eqqcolon C\left(T_{\delta}+T_{\rm loc}\right).$
Recall that $P_{\rm{loc}}(u)]^{ab}_{\ell k}=[{\mathcal{H}}(u)-{\mathcal{H}}^{\rm ref}]_{\ell k}^{ab}$ if $(\ell,k)\in B_{R_{\delta}}\times B_{R_{\delta}}$, and $P_{\rm loc}]_{\ell k}^{ab}=0$ if $\ell$ or $k\notin B_{R_{\delta}}$. Here, we simplify $R_{\delta}(=R_{1}+R_{2})$ as $R$ in (6.2), by first choosing $R_{1}$ and then taking $R_{2}$ sufficiently large. Using (6.16) together with Hölder’s inequality, we obtain
$T_{\delta}\coloneqq\sum_{mn}\left|(z-{\mathcal{H}}(u))^{-1}_{\ell m}\right|\left|P_{\delta}(u)_{mn}\right|\left|(z-{\mathcal{H}}^{\rm ref})^{-1}_{nk}\right|$
$\leq C_{\ell k}(u)\sum_{m,n\in\Lambda\setminus B_{R}}e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell m}}\Big(e^{-\frac{\sqrt{3}}{2}\gamma_{0}\mathfrak{m}r_{mn}}|u(m)-u(n)|\Big)e^{-\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref})r_{nk}}$
@@
$\leq C_{\ell k}(u)\Big(\sum_{m,n\in\Lambda\backslash B_{R}}e^{-2\Upsilon r_{mn}}|u(m)-u(n)|^{2}\Big)^{\frac{1}{2}}\Big(\sum_{m,n\in\Lambda\setminus B_{R}}e^{-\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell m}}e^{-(\sqrt{3}\gamma_{0}\mathfrak{m}-2\Upsilon)r_{mn}}e^{-2\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref})r_{nk}}\Big)^{\frac{1}{2}}$
@@
$\leq C_{\ell k}(u)\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R})}\Big(\sum_{m,n\in\Lambda\setminus B_{R}}e^{-\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell m}}e^{-\frac{\sqrt{3}}{2}\gamma_{0}\mathfrak{m}r_{mn}}e^{-2\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref})r_{nk}}\Big)^{\frac{1}{2}}$
(6.28) $\leq C_{\ell k}(u)\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R})}e^{-\eta_{-}r_{\ell k}},$
where we denote $\eta_{-}:=\frac{1}{4}\min\{\gamma_{\rm{CT}}(\tfrac{1}{2}\mathfrak{d}^{\mathrm{ref}}),\gamma_{0}\mathfrak{m}\}=c_{2}\frac{\gamma_{0}}{h_{0}}\min\{h_{0},\gamma_{0}^{d}\mathfrak{d}^{\mathrm{ref}}\}$ with $c_{2}>0$ depending only on $d,\mathfrak{m},M$ and $N_{\mathrm{b}}$. Furthermore, by combining the exponential sum estimate (C.2) with Combes–Thomas estimate (3.4) and its improved form (6.2), we obtain
$T_{\rm loc}$ $\coloneqq\sum_{mn}\left|(z-{\mathcal{H}}(u))^{-1}_{\ell m}\right|\big|P_{\rm loc}(u)_{mn}\big|\left|(z-{\mathcal{H}}^{\rm ref})^{-1}_{nk}\right|$
$\leq C_{\ell k}(u)\|Du\|_{\ell^{2}_{\Upsilon}(B_{R})}\left(\sum_{m\in B_{R}}e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell m}}\right)\left(\sum_{n\in B_{R}}e^{-\gamma_{\rm{CT}}(\mathfrak{d}^{\rm ref})r_{nk}}\right)$
(6.29) $\leq C_{\ell k}(u)\|Du\|_{\ell^{2}_{\Upsilon}(B_{R})}e^{\eta_{-}R}e^{-\eta_{-}(|\ell|+|k|)}$
From Lemma C, we note that the prefactors in (6.2) and (6.2) are both independent of $\eta_{-}$ as $\mathsf{gap}_{-}$ is small. Moreover, by choosing $\mathscr{C}$ satisfying $\mathrm{dist}(\mathscr{C},\sigma({\mathcal{H}}^{\mathrm{ref}}))\geq\frac{1}{2}\mathsf{gap}_{-}$, we can take $\mathsf{gap}_{-}$ in $\eta_{-}$. Substituting (6.2) and (6.2) into (6.2) draws the conclusion that
$\big|\rho(u)_{\ell k,ab}-\rho^{\mathrm{ref}}_{\ell k,ab}\big|$ $\leq C(T_{\delta}+T_{\rm loc})$
$\hskip-28.45274pt\leq C_{\ell k}(u)\Big(\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R})}+\|Du\|_{\ell^{2}_{\Upsilon}(B_{R})}e^{\eta_{-}R}e^{-\eta_{-}(|\ell|+|k|-r_{\ell k})}\Big)e^{-\eta_{-}r_{\ell k}}$
(6.30) $\hskip-28.45274pt\coloneqq\tilde{C}_{\ell k}(u)e^{-\eta_{-}r_{\ell k}}.$
Notice that $R=R_{1}+R_{2}$, and that the factor $C_{\ell k}(u)$ defined in (6.14) depends only on $R_{1}$. As $|\ell|+|k|-r_{\ell k}\to\infty$, we have $\tilde{C}_{\ell k}(u)\lesssim C\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\setminus B_{R})}$. Moreover, as $\|Du\|_{\ell^{2}_{\Upsilon}}\to 0$, taking the limit $R_{2}\to\infty$ yields $\tilde{C}_{\ell k}(u)\to 0$. Finally, substituting (6.2) into (6.26) completes the proof. ∎
6.3. Proof of the Theorem 3
Under assumption (Reg), we apply the mean value theorem to the bands $\varepsilon_{n}(\bm{\xi}+i\bm{\zeta}),~n\in\{N_{0},N_{0}+1\}$. Noting that $\nabla\varepsilon_{n}(\bm{\xi})$ is real, the first-order term $\operatorname{Re}(i\bm{\zeta}^{\top}\nabla\varepsilon_{n}(\bm{\xi}))$ vanishes, yielding: for all $\bm{\xi},\bm{\zeta}$ such that $\bm{\xi}+i\bm{\zeta}\in\mathbb{C}_{\delta_{*}}^{d}$, we have
(6.31) $\operatorname{Re}\varepsilon_{n}(\bm{\xi}+i\bm{\zeta})-\varepsilon_{n}(\bm{\xi})$ $=-\frac{1}{2}\bm{\zeta}^{\top}\nabla^{2}\varepsilon_{n}(\bm{\xi}^{\prime})\bm{\zeta},$
for some $\bm{\xi}^{\prime}\in[\bm{\xi},\bm{\xi}+i\bm{\zeta}]$. First, we estimate the norm of the Hamiltonian in the complex domain. For any $\bm{\xi}+i\bm{\zeta}\in\mathbb{C}_{\delta_{*}}^{d}$, we have